OpenStudy (anonymous):

14. Using elimination to solve the system 2x + y = 8 and x + y = 4, which of the following could be the first step? (Points : 2) 2x + y = 8--------> 2x + y = 8 (x + y = 4)(-1)---> -x - y = -4 2x + (4 - x) = 8 2x + y = x + y 3x + 2y = 12 x = 4 15. Solve the system x + y = 6 and x - y = 2. What is the value of x? (Points : 2) 6 2 4 0

OpenStudy (bahrom7893):

1)2x + (4 - x) = 8

OpenStudy (bahrom7893):

x+y=6 x-y=2 add the equations: 2x = 8 x = 4